# Whetten Maison Parcollet Dinarelli Estève Avignon UniversitéFrance Thales SIXFrance Samsung AI Center CambridgeUnited Kingdom Université Grenoble AlpesFrance

### v1

@@ Abstract
In Self-Supervised Learning (SSL), pre-training and evaluation is resource intensive. In the speech domain, current indicators of the quality of SSL models during pre-training, such as the loss, do not correlate well with downstream performance. Consequently, it is often difficult to gauge the final downstream performance in a cost efficient manner during pre-training. In this work, we propose unsupervised efficient methods that give insights into the quality of the pre-training of SSL speech models, namely, measuring the cluster quality and rank of the embeddings of the SSL model. Results show that measures of cluster quality and rank correlate better with downstream performance than the pre-training loss with only one hour of unlabeled audio, reducing the need for GPU hours and labeled data in SSL model evaluation.
@@ 1 Introduction
Self-Supervised Learning (SSL) is the process of training a model where the targets are derived directly from the input data itself, decreasing the need ofhuman-labeled data. These models are usually trained in two stages: pre-training, where the model is trained using unlabelled data, and subsequently, fine-tuning, where the model is trained for a specific task using labeled data [1]. In the speech processing domain, SSL models have performed well on a variety of tasks such as Automatic Speech Recognition (ASR) or Speaker Verification (SV) [1, 2].
@@
Although performant, SSL speech models are computationally and memory expensive, usually requiring anywhere from 8 - 128 high-end GPUs [3, 4], and up to tens of millions of hours of data [5]. While recent work has been done to reduce the computation and memory requirements necessary to train these models [3, 6, 7, 8], limited research has focused on the efficient evaluation of these models during pre-training. However, if it were possible to evaluate a model’s performance effectively during pre-training, the total number of models requiring full training could be significantly reduced, leading to significant energy and compute savings.
@@
While it is common practice to monitor the pre-training loss, it does not necessarily correlate well with the downstream performance [3]. Thus, the most reliable method of evaluation remains to simply test each model on the downstream tasks. This can be done by fine-tuning for a specific task, or by using benchmarks such as SUPERB or the Multi-Probe Speech Self-Supervision benchmark (MP3S) [2, 9], in which the pre-trained SSL weights are frozen, and only a light downstream model is trained. Though fine-tuning and benchmarking are feasible, these solutions require labeled data, additional GPU hours, and are impractical to conduct when pre-training many models.
@@
To overcome the need for labeled data and additional GPU hours, researchers in the field of computer vision proposed using the effective rank to evaluate the quality of SSL representations [10, 11]. This methodology has shown promising results in the speech domain on the tasks of Phoneme Recognition, Keyword Spotting, and Speaker Identification [12]. However, whether this methodology extends to other tasks such as ASR and SV, and can be used at early stages of pre-training to predict final downstream performance remain open questions. Furthermore, no work has been done to explore other measures such as clustering methods.
@@
In this work, we propose the following two efficient unsupervised methods as indicators of final downstream performance at early stages in pre-training: measuring the (i) clusters quality and the (i) rank of the embeddings outputted by SSL models. To test this methodology, we experiment with the BERT-based Speech pre-Training with Random-projection Quantizer (BEST-RQ) architecture [13]. We train 30 BEST-RQ models varying the batch size and the proportion of audio masked and focus on ASR and SV tasks.111code at LINK (withheld to preserve anonymity)
Results show that, for ASR, clustering measures and the rank of SSL embeddings correlate better with the final performance than the pre-training loss, even at early stages in training, having the potential to save thousands of GPU hours in future experiments.
@@ 2.1 Architecture of BEST-RQ
We use BEST-RQ due to its efficiency, performance, and open source implementation in the SpeechBrain library [13, 14, 15]. In BEST-RQ there are three main components: the quantizer, the acoustic feature extractor (AFE) consisting of two CNN layers, and the context encoder made up of a series of Conformer layers [16]. Given an audio, Mel filterbanks features are calculated and then given to the quantizer. The AFE also receives the Mel filterbanks features, except a random portion of Mel filterbanks are masked (replaced with noise). The output of the AFE is then passed to the context encoder.
Using a random frozen linear projectand a codebook lookup, the quantizer generates discrete pseudo-targets and context encoder is trained to predict these targets forthe masked portions. The loss is calculated by taking the cross entropy between the predicted targets and the pseudo-targets. Due to these training conditions, the context encoder must use the surrounding unmasked context in order to make accurate predictions on the masked portions [13].
@@ 2.2 Pre-training loss vs downstream performance
In this section, we show theoretically how certain hyper-parameters changes can affect the loss. Let us consider two SSL BEST-RQ models $M_{a}$ and $M_{b}$, with respective masking hyper-parameters $m_{a}$ and $m_{b}$, which represent the percentage of audio that is masked during pre-training. Let $m_{a}<m_{b}$ and where $m_{a}\approx 0$ meaning that almost none of the input is masked. Let both models use the same quantizer $Q$. Finally, let $\text{x}=[x_{1},x_{2},...x_{T}]$ be an input sequence of length $T$.
As the context encoder uses the global context to predict the targets of the masked sections, the larger the mask, the smaller amount of uncorrupted input available to the model. Thus, the model $M_{b}$, having a larger mask $m_{b}$, will have more uncertainty in its prediction, resulting in higher entropy, and consequently, a higher loss. However, having a higher loss is not always a negative factor in this scenario.
On the other hand, since $m_{a}\approx 0$, $M_{a}$ would be trained to mimic $Q$, essentially learning a single linear projection—a simple task for Conformers—resulting in a low loss. In other words, as $M_{a}$ is trained, the output of $M_{a}(x_{i})\approx Q(x_{i})$. Conversely, $M_{b}$, due to having a higher mask percentage $m_{b}$, needs to do more than just replicate $Q$, but instead learn statistical properties of natural language and audio signals. Thus, even though the loss of $M_{b}$ will be greater than the one of $M_{a}$, $M_{b}$ will learn important statistical properties which will lead to good initialization for downstream tasks. Other architectural and hyper-parameter changes, such as the size of the codebook, by similar logic, can have similar effects.
@@ 2.3 Using rank for unsupervised evaluation
The rank of a matrix is the number of linearly independent rows or columns and is a proxy for the information content of the matrix. Thus, the higher the rank, the less redundancy, the more information is contained by the matrix. The classical way to compute the rank of a matrix is to count the number of non-zero singular values. However, when working with real-world matrices which contain non-integer values, many, if not all of the singular values will be non-zero. This necessitates the use of a threshold. However, in practice this threshold is quite application-dependent and difficult to set. To solve this issue, researchers proposed using the effective rank (E-Rank) [11], which is a real-valued extension of the rank:
@@
where $A$ is a matrix of size $(N,M)$, and $(p_{i})_{i}$ are the normalized singular values. The effective rank, coined RankMe, was shown to be useful in the evaluation of self-supervised representations in computer vision [10].
@@ 3.1.1 Clustering Measures
Once a clustering model is trained, we use two metrics for evaluating the clustering quality, namely inertia [^2] Also called WCSS (Within-Cluster Sum of Squares) and Davies-Bouldin (DB) index [22]. For both of these measures, a lower value indicates a better clustering. For brevity, we will only define the DB index:
@@
where $k$ is the number of clusters, $s_{i}$ is the diameter of cluster $i$ and $d_{ij}$ is the distance between centroids of clusters $i$ and $j$. We experimented with other measures of cluster purity, but only report on inertia and the DB index.
@@ 3.2 Experimental settings
As previously noted, for our experiments we work with BEST-RQ due to its state-of-the-art performance, efficiency, and open source implementation. The efficiency is particularly important as we need to train many models to evaluate the methodology. We create various hyper-parameters configurations by varying the batch size from 15 to 120 minutes and the proportion of audio masked from 40% to 80%. We train all models for 200k steps on the 960 hours from the train-clean-100, train-clean-360 and train-other-500 splits in the LibriSpeech dataset [23]. We focus on two main checkpoints: 50k steps and the final 200k steps. While most SSL speech models are trained for longer [24, 25, 13], previous work has shown 200k steps to be sufficient for comparing SSL models [7, 14].
In our first downstream task, we benchmark, meaning the SSL model is frozen, the models on in-domain ASR data. For this we use the ContextNet LibriSpeech-100 task from the MP3S benchmark, where only a minimal ContextNet [26] probe is trained using the train-clean-100 subset of LibriSpeech. We evaluate using the WER on the dev-clean set of LibriSpeech with greedy decoding and no language model. As it common to fine-tune SSL models in practice, for the second task, we fine-tune on in-domain ASR data with the fine-tuning recipe in SpeechBrain. Similar to the first task, we evaluate the WER on the dev-clean set of LibriSpeech with greedy decoding and no language model. For our third task we benchmark (SSL model frozen), on out-of-domain ASR data. We use the LSTM CommonVoice 11.0 Basque ASR task from the MP3S benchmark. In this task a 2-layered BiLSTM probe is trained. The train and dev splits provided in CommonVoice 11.0 Basque dataset are use, and performance is evaluated using the WER on the development split. Lastly, in order to test these methods on a non-ASR task, we follow the MP3S benchmark and train SV models on VoxCeleb1. Performance is evaluated using the Error Rate on the development set.
@@
Table 2: Pearson Correlations between various unsupervised measures and future downstream performance. The unsupervised measures are calculated from the embeddings of SSL models at 50k steps pre-training and the correlation is measured with the final performance obtained by the same model after 200k steps of pre-training. Correlations for in-domain ASR remains high, but not for out-of-domain tasks. LS = LibriSpeech, CV = CommonVoice, VoxCel = VoxCeleb, (f.t.) = SSL Model is fine-tuned.
@@
Figure 1: Correlations between 4 selected measures (calculated on embeddings at 50k steps of pre-training) and the future WER on the LibriSpeech dev-clean set (at 200k steps of pre-training and 20 epochs of fine-tuning). Each point represents a model with different hyper-parameters (batch size and mask percentage). Fig. 1(a), show that some of the models with the lowest WER (y-axis) have a relatively high pre-training loss (x-axis) resulting in a low correlation, whereas unsupervised measures have a stronger correlation with WER.
@@ 4 Results
We first report on preliminary experiments describing how we decided the amount audio to use for the unsupervised evaluation methods and the number of epochs of the downstream task we perform. Then, we describe results for each downstream task.
@@ 4.1 Preliminary experiments
For generating the output embeddings, we randomly sample about 1 hour of audio from each dataset. This steamsfrom preliminary experiments in which we find that the global effective rank plateaus quickly does not increase much when using more than 1 hour of audio.
Another preliminary finding is a strong correlation between the WER early on in pre-training and the final WER when working with in-domain data. This means that in order to gauge performance with labeled data, one does not need to fully pre-train nor fully fine-tune when working with in-domain ASR. For an example, the Pearson correlations between WER from minimally trained models, 50k steps of pre-training and 5 epochs on the downstream task, and their corresponding final models, 200k steps of pre-training and 20 epochs on the downstream task, were 0.951 and 0.967 when freezing and when fine-tuning the SSL model, respectively. Thus, for the ASR experiments we only run the downstream for 5 epochs for the models at 50k checkpoints to save computation.
@@ 4.2 Automatic speech recognition
For in-domain ASR experiments, results were similar whether the SSL models were fine-tuned or frozen; both cluster and rank methods correlate better with WER than the pre-training loss baseline. Some of the highest correlations came from the GER ranging from -0.63 - 0.96. In contrast, the pre-training loss resulted in weak correlations ranging from -0.170 to -0.303 (first four columns of Table 1). Graphs comparing a select few of these measures with the WER are shown in Fig. 1.
When looking into the predictability of future SSL model performance we similarly found high correlations between rank and cluster measure at embedding at 50k steps of pre-training and final WER after 200k steps of pre-training and 20 epochs of fine-tuning. For example, the correlations between the GER and final WER were around -0.9, yet only around -0.3 for the pre-training loss (first two columns of Table 2).
When it comes to out-of-domain (CommonVoice) ASR experiments, correlations remain high at 50k steps, with the highest being -0.885 for the GER. However, the correlations at 200k are not strong, nor are the correlations between unsupervised measures at 50k and the final WER at 200k (Table 2). Examining why this might be the case, we observe that at 200k all the models perform similarly, with an absolute difference in WER of only 2.8 between the worst and best performance, suggesting that if model performance is very similar, it becomes hard to judge based on clustering or rank measures.
@@ 4.3 Speaker verification
With regards to out-of-domain SV, we observe correlations between unsupervised measures and the error rate after 20 epochs of fine-tuning. While they are not as strong as with ASR, the unsupervised measures still correlate better than the pre-training loss. Note that for 200k checkpoints, there is a strong negative correlation between the loss and the downstream performance of -0.598 (last column Table 1), indicating that the higher the loss, the lower the error rate which is not useful. We find that, like CommonVoice, the majority of models performed similarly (90% of the models exhibited error rates within a 1% range), leading to low correlations.

### v2

@@ Abstract
In Self-Supervised Learning (SSL), pre-training and evaluation are resource intensive. In the speech domain, current indicators of the quality of SSL models during pre-training, such as the loss, do not correlate well with downstream performance. Consequently, it is often difficult to gauge the final downstream performance in a cost efficient manner during pre-training. In this work, we propose unsupervised efficient methods that give insights into the pre-training quality of SSL speech models, namely, measuring the cluster quality and rank of the embeddings produced by the SSL model. Results show that measures of cluster quality and rank correlate better with downstream performance than the pre-training loss, reducing the need for GPU hours and labeled data in SSL model evaluation.
@@ 1 Introduction
Self-Supervised Learning (SSL) is the process of training a model where the targets are derived directly from the input data itself, decreasing the need of human-labeled data. SSL models are usually trained in two stages: pre-training, where the model is trained using unlabelled data, and subsequently, fine-tuning, where the model is trained for a specific task using labeled data [1]. In the speech processing domain, SSL models have performed well on a variety of tasks such as Automatic Speech Recognition (ASR) or Speaker Verification (SV) [1, 2].
@@
Although performant, SSL speech models are computationally and memory expensive, usually requiring anywhere from 8 - 128 high-end GPUs [3, 4], and up to tens of millions of hours of data [5]. While recent work has been done to reduce the computation and memory requirements necessary to train these models [3, 6, 7, 8], limited research has focused on the efficient evaluation of these models during pre-training. Nevertheless, if it were possible to evaluate a model’s performance effectively during pre-training, the total number of models requiring full training could be significantly reduced, leading to significant energy and compute savings.
@@
While it is common practice to monitor the SSL pre-training loss, it does not necessarily correlate well with the downstream performance [3]. Thus, the most reliable method of evaluation remains to simply test each model on the downstream tasks. This can be done by fine-tuning for a specific task, or by using benchmarks such as SUPERB or the Multi-Probe Speech Self-Supervision benchmark (MP3S) [2, 9], in which the pre-trained SSL weights are frozen, and only a light downstream model, sometimes called a probe, is trained. Though these evaluation methods are feasible, they require labeled data and additional GPU hours.
@@
To overcome the need for labeled data and additional GPU hours, researchers in the field of computer vision proposed using the effective rank to evaluate the quality of SSL representations [10, 11]. This methodology has shown promising results in the speech domain on the tasks of Phoneme Recognition, Keyword Spotting, and Speaker Identification [12]. However, whether this methodology extends to other tasks such as ASR and SV remains an open question. Additionally, no work has explored other measures like clustering methods. Furthermore, whether this approach can be used early in pre-training to predict final downstream performance remains unclear.
@@
To address these gaps, we propose two compute- and data-efficient unsupervised methods as indicators of final downstream performance at early stages of pre-training: measuring the clusters quality and the rank of the embeddings outputted by SSL models. To test these methods, we experiment with the BERT-based Speech pre-Training with Random-projection Quantizer (BEST-RQ) architecture [13], training 30 BEST-RQ models varying the batch size and the proportion of audio masked and focus on ASR and SV tasks.111code at https://github.com/whettenr/ssl-speech-evaluation Results show that, for ASR, clustering measures and the rank of SSL embeddings correlate better with the final performance than the pre-training loss, even at early stages in training, having the potential to save thousands of GPU hours in future experiments.
@@ 2.1 Architecture of BEST-RQ
We use BEST-RQ due to its efficiency, performance, and open source implementation in the SpeechBrain library [14, 15]. In BEST-RQ there are three main components: the quantizer, the acoustic feature extractor (AFE) consisting of two CNN layers, and the context encoder made up of a series of Conformer layers [16]. Given an audio, Mel filterbanks features are calculated and then given to the quantizer. The AFE also receives the Mel filterbanks features, except a random portion of Mel filterbanks are masked (replaced with random noise). The output of the AFE is then passed to the context encoder.
Using a random frozen linear projection and a codebook lookup, the quantizer generates discrete pseudo-targets and the context encoder is trained to predict these targets for the masked portions. The loss is calculated by taking the cross entropy between the predicted targets and the pseudo-targets. Due to these training conditions, the context encoder must use the surrounding unmasked context in order to make accurate predictions on the masked portions [13].
@@ 2.2 Pre-training loss vs downstream performance
In this section, we show theoretically how certain hyper-parameters changes can affect the loss. Let us consider two SSL BEST-RQ models $M_{a}$ and $M_{b}$, with respective masking hyper-parameters $m_{a}$ and $m_{b}$, which represent the percentage of audio that is masked during pre-training. Let $m_{a}<m_{b}$ and $m_{a}\approx 0$ meaning almost none of the input is masked. Let both models use the same quantizer $Q$. Finally, let $\text{x}=[x_{1},x_{2},...x_{T}]$ be an input sequence of vectors of length $T$.
As the context encoder uses the global context to predict the targets of the masked sections, the larger the mask, the smaller amount of uncorrupted input available to the model. Thus, the model $M_{b}$, having a larger mask $m_{b}$, will have more uncertainty in its prediction, resulting in higher entropy, and consequently, a higher loss. However, as will be explained subsequently, having a higher loss is not always a negative factor in this scenario.
Since $m_{a}$ is very small, $M_{a}$ would be trained to mimic $Q$, essentially learning a single linear projection—a simple task for Conformers—resulting in a low loss. In other words, as $M_{a}$ is trained, the output of $M_{a}(x_{i})\approx Q(x_{i})$. Conversely, because $M_{b}$ has a higher mask percentage, it does more than simply replicate $Q$. Instead, it uses the unmasked context to make accurate predictions of $Q$, compelling $M_{b}$ to learn the statistical properties of speech and audio. As a result, while $M_{b}$ incurs a higher loss than $M_{a}$, it learns important statistical properties that contribute to a strong initialization for downstream tasks.
@@ 2.3 Using rank for unsupervised evaluation
The rank of a matrix is the number of linearly independent rows or columns and is a proxy for the information content of the matrix. Thus, the higher the rank, the less redundancy, the more information is contained in the matrix. The classical way to compute the rank of a matrix is to count the number of non-zero singular values. However, when working with real-world matrices which contain non-integer values, many, if not all of the singular values will be non-zero. This necessitates the use of a threshold. However, in practice this threshold is quite application-dependent and difficult to set. To solve this issue, researchers proposed using the effective rank (E-Rank) [11], which is a real-valued extension of the rank:
@@
where $A$ is a matrix of size $(N,M)$, and $p_{i}$ are the normalized singular values. The effective rank, coined RankMe, was shown to be useful in the evaluation of self-supervised representations in computer vision [10].
@@ 3.1.1 Clustering measures
Once a clustering model is trained, we use two metrics for evaluating the clustering quality: inertia, also called the Within-Cluster Sum of Squares (WCSS), and the Davies-Bouldin (DB) index [22]. The WCSS measures the average intra-cluster distance, or the mean distance between each data point and its assigned cluster centroid. This can be expressed by the following:
$\textrm{WCSS}=\sum_{k=1}^{K}\sum_{i\in C_{k}}||x_{i}-\mu_{k}||^{2},$ (2)
where $K$ is the number of custers, $C_{k}$ is the set of data points assigned to cluster $k$, $x_{i}$ is a single data point and $\mu_{k}$ is the centroid of cluster $k$. In contrast to the WCSS, the DB index takes into the intra- and inter-cluster distance and is calulated by the following:
@@
where $K$ is the number of clusters, $s_{i}$ is the diameter of cluster $i$ and $d_{ij}$ is the distance between centroids of clusters $i$ and $j$.
@@ 3.2 Experimental settings
As previously noted, for our experiments we work with BEST-RQ due to its state-of-the-art performance, efficiency, and open source implementation. The efficiency is particularly important as we need to train many models to evaluate the methodology. We create various hyper-parameters configurations by varying the batch size from 15 to 120 minutes and the proportion of audio masked from 40% to 80%. We train all models for 200k steps on the 960 hours set of LibriSpeech [23]. We focus on two main checkpoints: 50k steps and 200k steps. While most SSL speech models are trained for longer [24, 25, 13], previous work has shown 200k steps to be sufficient for evaluation [7, 14].
For our first task, we evaluate the frozen SSL models on in-domain ASR using the ContextNet LibriSpeech-100 task from the MP3S benchmark [9]. In this task, a minimal ContextNet [26] probe is trained on the train-clean-100 subset of LibriSpeech, and performance is measured by word error rate (WER) on the dev-clean set using greedy decoding without a language model. Since fine-tuning is common in practice, we also fine-tune the SSL models on in-domain ASR data using the SpeechBrain fine-tuning recipe. Evaluation follows the same setup as the frozen model, measuring WER on the dev-clean set with greedy decoding and no language model. For our third task, we evaluate the frozen SSL models on out-of-domain ASR using the LSTM CommonVoice 11.0 Basque ASR task from MP3S. For this task, a 2-layer BiLSTM probe is trained on the provided train splits. Evaluation is done using the WER on the provided development set. Lastly, to assess performance on a non-ASR task, we follow the MP3S benchmark to evaluate the frozen SSL models on Speaker Verification using VoxCeleb1 [27]. In this task, a ECAPA-TDNN [28] probe is trained and models are evaluated using error rate on the development set.
@@
Table 2: Pearson Correlations between various unsupervised measures and future downstream performance. The unsupervised measures are calculated from the embeddings of SSL models at 50k pre-training steps and the correlation is measured with their respective final performances at 200k pre-training steps. Correlations for in-domain ASR remains high, but not for out-of-domain tasks. LS = LibriSpeech, CV = CommonVoice, VoxCel = VoxCeleb, (f.t.) = SSL Model is fine-tuned.
@@
Figure 1: Correlations between 4 selected measures (calculated on embeddings at 50k pre-training steps) and the future WER on the LibriSpeech dev-clean set (at 200k pre-training steps and 20 epochs fine-tuning). Each point represents a model with different hyper-parameters (batch size and mask percentage). Fig. 1(a), show that some of the models with the lowest WER (y-axis) have a relatively high pre-training loss (x-axis) resulting in a low correlation, whereas unsupervised measures have a stronger correlation with WER.
@@ 4 Results
We first report on preliminary experiments that informed our choices on the amount of audio to use for the unsupervised evaluation methods and the number of epochs for the downstream task. Then we present results for each task.
@@ 4.1 Preliminary experiments
For generating the output embeddings, we randomly sample about 1 hour of audio from each dataset. This stems from preliminary experiments in which we find that the global effective rank plateaus quickly, increasing minimally when using more than 1 hour of audio. Computing all embeddings and metrics for a single layer of one model takes about 20 minutes on CPUs, with the clustering accounting for most of the time.
Another preliminary finding is a strong correlation between the WER early on in pre-training and the final WER when working with in-domain data. For example, the Pearson correlations between WER from minimally trained models (50k pre-training steps and 5 downstream epochs) and their final counterparts (200k pre-training steps, 20 downstream epochs) were 0.951 and 0.967 when freezing and fine-tuning the SSL model, respectively. This means that in order to gauge performance with labeled data, one does not need to fully pre-train nor fully fine-tune when working with in-domain ASR. Thus, for the ASR experiments we only run the downstream for 5 epochs for the models at 50k checkpoints to save computation.
@@ 4.2 Automatic speech recognition
For in-domain ASR experiments, results were similar whether the SSL models were fine-tuned or frozen; both cluster and rank methods correlate with the WER better than the pre-training loss baseline (see first four columns of Table 1). Some of the highest correlations came from the GER ranging from -0.63 to -0.96, meaning the higher the rank the better the performance (lower WER). In contrast, the pre-training did not correlated strongly with WER, with correlations ranging from -0.170 to -0.303. Graphs comparing a select few of these measures with the WER are shown in Fig. 1.
Similarly, we find high correlations between clustering and rank measures at 50k steps of pre-training and final WER after 200k steps of pre-training and 20 epochs of fine-tuning. This sugests that, without any fine-tuning, we can guage the finial performance very early on in pre-training for in-domain ASR. For example, the correlations between the GER and final WER were around -0.9, yet only around -0.3 for the pre-training loss (first two columns of Table 2).
When it comes to out-of-domain (CommonVoice) ASR experiments, correlations between unsupervised measure and WER remain high at 50k steps, with the highest being -0.885 for the GER. However, the correlations at 200k are not strong, nor are the correlations between unsupervised measures at 50k and the final WER at 200k (Table 2). Examining why this might be the case, we observe that at 200k all the models perform similarly, with an absolute difference in WER of only 2.8 between the worst and best performance, suggesting that if model performance is very similar, it becomes hard to judge based on clustering or rank measures.
@@ 4.3 Speaker verification
With regards to Speaker Verification, we observe correlations between unsupervised measures and the error rate after 20 epochs of fine-tuning. While they are not as strong as with ASR, the unsupervised measures still correlate better than the pre-training loss. Note that for 200k checkpoints, there is a strong negative correlation between the loss and the downstream performance of -0.598 (last column Table 1), indicating that the higher the loss, the lower the error rate which is not useful. We find that, like CommonVoice, the majority of models performed similarly (90% of the models exhibited error rates within a 1% range), leading to low correlations.